A manufacturer wishes to fabricate a step index optical fibre from a material such that relative core-cladding index difference is 0.01. The refractive index of cladding is 1.3. In other case, keeping the material of the core same, the cladding material is replaced by another medium, such that core-cladding index difference is now 0.02. Calculate the ratio of light gathering ability of the optical fibre for the two cases with refractive index of outer medium to be 1. Compare the two cables for their applications in communication. Suggest the measures to be done if graded index fibre is to be fabricated in both the cases and how it is going to change the properties and applications of cable.


1
Expert's answer
2021-11-24T12:00:50-0500

"\Delta=\frac{n_2'}{n_1'\sqrt{n_2^2-n_1^2}}=75."